OpenStudy (anonymous):

Type this equation in standard form. 3x = 2y + 7 @ivancsc1996

OpenStudy (ivancsc1996):

Ok, so standard form I am guessing is function form, that is yo write y as a unction of x. For that you just need to solve for y. Substract by 7 the whole thing and then divide by 2 and thats it.\[3x=2y+7 \rightarrow y=\frac{ 3x-7 }{ 2 }\]

OpenStudy (anonymous):

how do u do that

OpenStudy (anonymous):

i got the answer its 3x-2y=7

OpenStudy (ivancsc1996):

ok, so you ffirst substarct 7 from both sides of the equation.\[3x-7=2y+7-7 \rightarrow 3x-7=2y\]Then you divide by two on both sides:\[\frac{ 3x-7 }{ 2 }=\frac{ 2y}{ 2 } \rightarrow \frac{ 3x-7 }{ 2 }=y\]Any doubts? You can do does additions and divisions, since you are doing it to BOTH sides, therefore keeping the equation balanced. Remember, whatever you do to one side, you need to do to the other.